Find the angle between the vectors.
1. v=(-3,4) u=(8,5)
2. (-3,8), (-1,-9) this one did not come with u and v.

Found 2 solutions by Alan3354, ikleyn:
Answer by Alan3354(69443)

Find the angle between the vectors.
1. v=(-3,4) u=(8,5)
Dot product = |v|*|u|*cos(A)
-3*8 + 4*5 = 5*sqrt(89)*cos(A)
cos(A) = -4/(5sqrt(89)) =~ -0.08479983
Angle =~ 94.8645 degs
================
2. (-3,8), (-1,-9) this one did not come with u and v.
Same problem, different numbers.

Answer by ikleyn(52780)

.
See relevant lessons on dot-product
    - Introduction to dot-product
    - Formula for Dot-product of vectors in a plane via the vectors components
    - Dot-product of vectors in a coordinate plane and the angle between two vectors
    - Perpendicular vectors in a coordinate plane
    - Solved problems on Dot-product of vectors and the angle between two vectors
    - Properties of Dot-product of vectors in a coordinate plane
in this site.

Also,  you have this free of charge online textbook in ALGEBRA-II in this site
    - ALGEBRA-II - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic
"Dot-product for vectors in a coordinate plane".